A teacher wants to estimate the mean time (in minutes) that students take to go from one classroom to the next. His research assistant uses the sample time of 40 students to report the confidence interval as [7.40, 8.60].


a. Find the sample mean time used to compute the confidence interval. (Round intermediate calculations to 4 decimal places and final answer to the nearest whole number.)  

Sample Mean -

b. Determine the confidence level if the sample standard deviation used for the interval is 1.876. (Round intermediate calculations to at least 4 decimal places. Round "t" value to 3 decimal places and final answer to the nearest whole number.)

Confidence Level % -
Answer #1

sample size,n = 40

A)

sample mean,x = (7.40+8.60)/2

= 8

B)

standard deviation,s = 1.872

margin of error,E = Za/2 * (s/sqrt(n))

E = (8.60-7.40)/2

E = 0.60

E = Za/2 * (s/sqrt(n)) = 0.60

Za/2 =0.0212

a/2 = 0.494

a = 0.9576

= 95.76% confidence level